You have $68.00 and you must find the percent that you subtract from that number to get $35.00.
Answer by nerdybill(7384)

You have $68.00 and you must find the percent that you subtract from that number to get $35.00.
.
Let x = percent
then
68 - 68x = 35
-68x = -33
x = (-33)/(-68)
x = 0.485
answer: 48.5%